Suppose that a random variable X follows a binomial distribution. Find the bias of each of the following estimators. Be sure to follow your computations enough to determine whether the estimator is biased or unbiased.

a. p̂ = (X+1)/(n+2)

b. p̂ = X2/n2

c. p̂ = X/n
Answer #1

Here,

X ~ Binomial (n, p)

So, E(X) = n*p and Var(X) = n*p*(1-p)

a) n2 n2 7l

So, hat{p} is a biased estimator for p.

b) XE(X2) Var(X)E(X)p(1 - p) +p

Rightarrow E(hat{p}) = rac{np(1-p) + p^2}{n^2}

Rightarrow E(hat{p}) = rac{p(1-p)}{n} + rac{p^2}{n^2}

Rightarrow E(hat{p}) eq p

So, hat{p} is a biased estimator for p.

c) E(p) = E(X) = E(X)--n * p

So, hat{p} is a unbiased estimator for p.